# Eisenstein degeneration of Euler systems

### v1

@@ 5.5. Meromorphic Eichler–Shimura
If $L_{p}^{\prime}(f_{\beta},\chi)$, there exists an integer $n\geq 0$ and a unique isomorphism of $\mathcal{O}(U)$-modules
$\eta_{{\mathbf{f}}}\colon\mathbf{D}_{\mathrm{cris}}(\mathcal{F}^{-}D({\mathbf{f}})^{*})\cong X^{-n}\mathcal{O}(U)$
@@
Let $n\geqslant 0$ be the order of vanishing of $d_{{\mathbf{f}}}$ at $k=r$ and suppose that Assumption 1.1 holds and $L_{p}^{\prime}(f_{\beta},\chi)\neq 0$. Then $X^{-n}\kappa({\mathbf{f}})$ is well-defined and non-zero modulo $X$; and this leading term projects into the quotient $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}))$. Its image under the Perrin-Riou regulator is given by
$d^{*}_{{\mathbf{f}}}(r)\cdot\mathcal{E}_{N}\cdot\frac{L_{p}^{\prime}(f_{\beta},\chi)}{L_{p}^{\prime}(\operatorname{Ad}f_{\beta})}\cdot L_{p}(E_{r+2}(\psi,\tau),{\mathbf{j}}+1),$
@@
This directly follows from the previous discussion and the definition of $d^{*}_{{\mathbf{f}}}(r)$. ∎
@@
$L_{p}(f_{\beta},{\mathbf{j}}+1)=\lambda\cdot\log_{p}^{[r+1]}\cdot L_{p}(\psi,{\mathbf{j}}+1)\cdot L_{p}(\tau,{\mathbf{j}}-r).$
@@ 5.6. Comparison with circular units
We have that $L_{p}^{\prime}(\operatorname{Ad}{\mathbf{f}})(r)\neq 0$. If Assumption 5.15 holds true, then we further have that $L_{p}(f,\chi)(k,k/2)$ vanishes at $k=r$.
@@
The vanishing of $L_{p}(f,\chi)(k,k/2)$ follows by Theorem 5.6. ∎
@@
For $n=0$, it is clear that the class $\hat{\kappa}^{*}(f_{\beta})$ is divisible by $\log^{[r+1]}$; for $n>0$ this is less obvious, but it follows from the proof of the next theorem. For the statement, we implicitly assume that $L_{p}(\operatorname{Ad}{\mathbf{f}})$ has a simple zero at $k=r$; if not, we must understand it by multiplying both sides of the equality by $L_{p}^{\prime}(\operatorname{Ad}{\mathbf{f}})$, thus having a zero in the right hand side.
@@
Assume that $r$ is even and that both $\psi$ and $\tau$ are even. Under Assumption 1.1 and Assumption 1.2, and if we further impose that $L_{p}^{\prime}(f_{\beta},\chi)\neq 0$, the following equality holds in $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}})))$:
$\kappa^{*}(f_{\beta})=C\cdot\lambda\cdot\frac{\mathcal{E}_{N}\cdot\log^{[r+1]}\cdot L_{p}^{\prime}(f_{\beta},\chi)\cdot L_{p}(\psi,{\mathbf{j}}+1)}{L_{p}^{\prime}(\operatorname{Ad}f_{\beta})}c(\tau)(r),$
where $C$ is an explicit non-zero constant which does not depend on $j$.
@@
As it occurred in the Beilinson–Flach case, there is a logarithmic contribution in the formula. However, in [LRS24] that contribution came from the class itself, since we had been able to argue, in terms of the geometry, that it was divisible by $\log^{[r+1]}$. Here, the term arises in the factorization formula of the $p$-adic $L$-function.
As we have already discussed, the terms $L_{p}^{\prime}(f_{\beta},\chi)$ and $L_{p}^{\prime}(\operatorname{Ad}f_{\beta})$ do not depend on the cyclotomic variable, so we may drop them from the statement.
@@
Assume that $r$ is even and that both $\psi$ and $\tau$ are even. Under Assumption 1.1 and Assumption 1.2, and if we further impose that $L_{p}^{\prime}(f_{\beta},\chi)\neq 0$, the following equality holds in $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}})))$:
$\kappa^{*}(f_{\beta})=\tilde{C}\cdot\lambda\cdot E_{N}\cdot\log^{[r+1]}\cdot L_{p}(\psi,{\mathbf{j}}+1)\cdot c(\tau)(r),$
where $\tilde{C}$ is an explicit non-zero constant which does not depend on $j$.
@@ 6.2. Vanishing of the Kato class
$\operatorname{Col}_{b_{f}^{+}}=\left\langle\mathcal{L}^{\mathrm{PR}}_{\mathcal{F}^{+}}(-),b_{f}^{+}\right\rangle:H^{1}_{\mathrm{Iw}}(\mathbf{Q}_{p,\infty},\mathcal{F}^{+}D(f)^{*})\to\mathcal{O}(\mathcal{W}).$
@@
$\operatorname{Col}_{b_{f}^{+}}(\hat{\kappa}(f_{\beta}))\cdot L_{p}^{\prime}(\operatorname{Ad}f_{\beta})=c({\mathbf{k}})\cdot L_{p}(f_{\beta},\chi,r/2+1)\cdot L_{p}^{\prime}(f_{\beta},{\mathbf{j}}_{+}1),$
where $c({\mathbf{k}})$ is a meromorphic function on $U$ alone, regular and non-vanishing at all integer weights $k\geqslant-1$ except possibly at $r$ itself, and regular at $k=r$.
@@
When the character $\tau$ is odd and $\tau(p)\neq 1$, the space $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))$ is zero and the Beilinson–Kato $\hat{\kappa}(f_{\beta})\in H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))$ class vanishes.
@@ 6.3. The case of odd weight
The following table summarizes the four possible cases according to parity, assuming the absence of exceptional zeros. The BD column indicates with a check-mark whether the factorization formula of [BD15] (as a product of two Kubota–Leopoldt $p$-adic $L$-function) holds for $L_{p}({\mathbf{f}})$. The columns $\kappa(f_{\beta})$ and $\kappa(f_{\beta})_{2}$ contain a check-mark if the corresponding elements $\kappa(f_{\beta})$ and $\kappa(f_{\beta})_{2}$ do not vanish for reasons related to parity.
@@
It remains to consider the case where $\tau$ is odd and $\tau(p)=1$. From the point of view of Selmer groups, the exceptional vanishing of the $p$-adic $L$-function $L_{p}(\tau,s)$ when $\tau(p)=1$ corresponds to the existence of an extended Selmer group, which in this case is just a $p$-unit. For simplicity, consider the case of $r=j=0$, and write $\kappa^{*}(f_{\beta})(1)$ for the specialization of the Iwasawa class $\kappa^{*}(f_{\beta})$ corresponding to $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1))$. As a piece of notation, let $H$ stand for the number field cut out by $\tau$ and let $O_{H}^{\times}[1/p][\tau]$ denote the space of $p$-units in $H$ where the Galois action is via $\tau$. This space has a canonical generator, that we denote by $v_{\tau}$. This suggests the following question regarding how to interpret the possible torsion of Iwasawa cohomology.
@@
7. Beilinson–Flach classes and the triple product case
@@
Let ${\mathbf{f}}$ be a Coleman family passing through a critical Eisenstein series and let ${\mathbf{g}}$ be a (ordinary) Hida family. When studying the comparison between Beilinson–Flach classes and Beilinson–Kato classes in [LRS24] we had used the Hida–Rankin $p$-adic $L$-function $L_{p}^{g}({\mathbf{f}},{\mathbf{g}})$, and not the one corresponding to the situation where the critical Eisenstein family is dominant. In the framework explored in this article, the only option was precisely considering that situation, since we just had one family. Hence, it is natural to extend these ideas to the Hida–Rankin setting and ask the following question (which also admits an analogue in the setting of triple products).
@@
In the case of $L_{p}^{f}(E_{r+2},{\mathbf{g}})$, we expect to get some logarithmic contribution, as in [BD15]. The same question arises in the setting of triple products, where one would expect a similar situation for the triple product $p$-adic $L$-function of Andreatta–Iovita [AI21] when the dominant family passes through a critial Eisenstein point.
@@
Therefore, one may wonder if the specialization of the three-variable $p$-adic $L$-function $L_{p}^{f}({\mathbf{f}},{\mathbf{g}})$ at $f_{\beta}$ is zero. A priori, it is not correct to say that the $p$-adic $L$-function is the image under the Perrin-Riou map of a trivial projection of the Beilinson–Flach class, since in the reciprocity law we may have a zero in the term $d_{{\mathbf{f}}}({\mathbf{k}})$ which arises in the leading term argument.

### v3

@@ 5.5. Meromorphic Eichler–Shimura
Under the running assumptions, there exists an integer $n\geq 0$ and a unique isomorphism of $A_{U}$-modules
$\eta_{{\mathbf{f}}}\colon\mathbf{D}_{\mathrm{cris}}(\mathcal{F}^{-}D({\mathbf{f}})^{*})\cong X^{-n}A_{U}$
@@
Let $n\geqslant 0$ be the order of vanishing of $d_{{\mathbf{f}}}$ at $k=r$ and suppose that Assumption 1.1 and 1.2 hold. Then $X^{-n}\kappa({\mathbf{f}})$ is well defined and non-zero modulo $X$; and this leading term projects into the quotient $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}))$. Its image under the Perrin-Riou regulator is given by a multiple of
$\left(\frac{d^{*}_{{\mathbf{f}}}(r)\cdot\mathcal{E}_{N}\cdot L_{p}^{\prime}(f_{\beta},\chi)}{L_{p}^{\prime}(\operatorname{Ad}f_{\beta})}\right)\cdot L_{p}(E_{r+2}(\psi,\tau),{\mathbf{j}}+1),$
@@
This follows directly from the previous discussion and the definition of $d^{*}_{{\mathbf{f}}}(r)$. ∎
@@
$L_{p}(f_{\beta},{\mathbf{j}}+1)=\lambda\cdot\log^{[r+1]}\cdot L_{p}(\psi,{\mathbf{j}}+1)\cdot L_{p}(\tau,{\mathbf{j}}-r).$
@@ 5.6. Comparison with circular units
The following result justifies why we have considered $L_{p}^{\prime}(\operatorname{Ad}f_{\beta})$ and $L_{p}^{\prime}(f_{\beta},\chi)$.
@@
If Assumption 5.16 holds true, then $L_{p}(\operatorname{Ad}f_{\beta})=0$ and the (one-variable) $p$-adic $L$-function $L_{p}({\mathbf{f}},\chi)(k,k/2)$ vanishes at $k=r$.
@@
The vanishing of $L_{p}(\operatorname{Ad}f_{\beta})$ follows from the aforementioned results of [Bel21, §9]. Further, $L_{p}({\mathbf{f}},\chi)(k,k/2)=0$ as a consequence of Theorem 5.6. ∎
@@
Assume that $r$ is even and that both $\psi$ and $\tau$ are even. Under Assumption 1.1 and Assumption 1.2, the following equality holds in $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}})))$:
$\kappa^{*}(f_{\beta})=\left(\frac{C\cdot\lambda\cdot\mathcal{E}_{N}\cdot L_{p}^{\prime}(f_{\beta},\chi)}{L_{p}^{\prime}(\operatorname{Ad}f_{\beta})}\right)\cdot L_{p}(\psi,{\mathbf{j}}+1)c(\tau)(r),$
where $C$ is an explicit constant which does not depend on $j$.
@@
Contrarily to which occurs in the Beilinson–Flach case, there is not a logarithmic contribution in the formula. In [LRS24] that contribution came from the class itself, since we had been able to argue, in geometric terms, that it was divisible by $\log^{[r+1]}$. Here, although a term of that kind arises in the factorization formula of the $p$-adic $L$-function, it is canceled out by the Coleman map.
As we have already discussed, the terms $L_{p}^{\prime}(f_{\beta},\chi)$ and $L_{p}^{\prime}(\operatorname{Ad}f_{\beta})$ do not depend on the cyclotomic variable, so we may drop them from the statement. Roughly speaking, we can say that $\kappa^{*}(f_{\beta})$ agrees with $c(\tau)(r)$ multiplied by the Kubota–Leopoldt $p$-adic $L$-function of $\psi$.
@@
Assume that $r$ is even and that both $\psi$ and $\tau$ are even. Under Assumption 1.1 and Assumption 1.2, the following equality holds in $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}))$:
$\kappa^{*}(f_{\beta})=\tilde{C}\cdot\lambda\cdot\mathcal{E}_{N}\cdot L_{p}(\psi,{\mathbf{j}}+1)\cdot c(\tau)(r),$
where $\tilde{C}$ is an explicit constant which does not depend on $j$.
@@ 6.2. Vanishing of the Kato class
$\operatorname{Col}_{b_{f}^{+}}=\left\langle\mathcal{L}^{\mathrm{PR}}_{\mathcal{F}^{+}}(-),b_{f}^{+}\right\rangle:H^{1}_{\mathrm{Iw}}(\mathbf{Q}_{p,\infty},\mathcal{F}^{+}D(f)^{*})\to\mathcal{H}_{\Gamma}.$
@@
$\operatorname{Col}_{b_{f}^{+}}(\kappa^{*}(f_{\beta}))\cdot L_{p}^{\prime}(\operatorname{Ad}f_{\beta})=c({\mathbf{k}})\cdot L_{p}(f_{\beta},\chi,r/2+1)\cdot L_{p}^{\prime}(f_{\beta},{\mathbf{j}}_{+}1),$
where $c({\mathbf{k}})$ is a meromorphic function on $U$ alone, which is identically zero or which is regular and non-vanishing at all integer weights $k\geqslant-1$ except possibly at $r$ itself, and regular at $k=r$.
@@
When the character $\tau$ is odd and $\tau(p)\neq 1$, the space $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1+r)\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))$ is zero and the Beilinson–Kato $\kappa^{*}(f_{\beta})\in H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})\otimes\mathcal{H}_{\Gamma}(-{\mathbf{j}}_{+}))$ class vanishes.
@@ 6.3. The case of odd weight
Table 1 summarizes the four possible cases according to parity, assuming the absence of exceptional zeros. The BD column indicates with a check-mark whether the factorization formula of [BD15] (as a product of two Kubota–Leopoldt $p$-adic $L$-function) holds for $L_{p}({\mathbf{f}})$. The columns $\kappa(f_{\beta})$ and $\kappa(f_{\beta})_{2}$ contain a check-mark if the corresponding elements $\kappa(f_{\beta})$ and $\kappa(f_{\beta})_{2}$ do not vanish for reasons related to parity.
@@
It remains to consider the case where $\tau$ is odd and $\tau(p)=1$. From the point of view of Selmer groups, the exceptional vanishing of the $p$-adic $L$-function $L_{p}(\tau,s)$ when $\tau(p)=1$ corresponds to the existence of an extended Selmer group, which in this case is just a $p$-unit. For simplicity, consider the case of $r=j=0$, and write $\kappa^{*}(f_{\beta})(1)$ for the specialization of the Iwasawa class $\kappa^{*}(f_{\beta})$ corresponding to $H^{1}(\mathbf{Q},\mathbf{Q}_{p}(\tau^{-1})(1))$. As a piece of notation, let $H$ stand for the number field cut out by $\tau$ and let $\mathcal{O}_{H}^{\times}[1/p][\tau]$ denote the space of $p$-units in $H$ where the Galois action is via $\tau$. This space has a canonical generator, that we denote by $v_{\tau}$. This suggests the following question regarding how to interpret the possible torsion of Iwasawa cohomology.
@@
7. Extensions of this work: Beilinson–Flach classes and the triple product case
@@
Let $\mathbf{f}$} be a Coleman family passing through a critical Eisenstein series and let $\mathbf{g}$} be a (ordinary) Hida family. When studying the comparison between Beilinson–Flach classes and Beilinson–Kato classes in [LRS24] we had used the Hida–Rankin $p$-adic $L$-function $L_{p}^{g}({\mathbf{f}},{\mathbf{g}})$, and not the one corresponding to the situation where the critical Eisenstein family is dominant, that we would denote by $L_{p}^{f}({\mathbf{f}},{\mathbf{g}})$. In the framework explored in this article, the only option was precisely considering that situation, since we just had one family. Hence, it is natural to extend these ideas to the Hida–Rankin setting and ask the following question (which also admits an analogue in the setting of triple products).
@@
In the case of $L_{p}^{f}(E_{r+2},{\mathbf{g}})$, we expect to get some logarithmic contribution, as in [BD15], together with a logaritmic term. The same question arises in the setting of triple products, where one would expect a similar situation for the triple product $p$-adic $L$-function of Andreatta–Iovita [AI21] when the dominant family passes through a critial Eisenstein point.
@@
Therefore, one may wonder if the specialization of the three-variable $p$-adic $L$-function $L_{p}^{f}({\mathbf{f}},{\mathbf{g}})$ at $f_{\beta}$ is zero. A priori, it is not correct to say that the $p$-adic $L$-function is the image under the Perrin-Riou map of a trivial projection of the Beilinson–Flach class, since in the reciprocity law we may have a zero in the term $d_{{\mathbf{f}}}({\mathbf{k}})$ which arises in the leading term argument and one needs to be more cautious. Extending the techniques we have developed in this work, we are able to get a satisfactory answer of 7.1 in the second part of this paper, and also to understand better this issue.